# What makes unlearning hard and what to do about it

### v1

@@ A.6 Detailed results
The first factor affecting unlearning difficulty, as discussed in Section 4.1, is the degree of entanglement between the forget and retain sets in the embedding space. Specifically, the more entangled these sets are, the harder it becomes to unlearn. We present detailed results on how this factor affects unlearning difficulty in Table 4, 5, 6, and Figure 7. Table 4 displays the ToW results for different ES partitions for both CIFAR-10 and CIFAR-100. Comprehensive data, including forget, retain, test accuracy, and MIA performance, are provided in Figure 7, as well as in Table 5 for CIFAR-10 and Table 6 for CIFAR-100. All the results are averaged over 3 runs for each algorithm (6 runs for relabelling-based algorithms due to their higher variance), along with 95% confidence intervals.
@@
(a) CIFAR-10 with ResNet-18.
@@
(b) CIFAR-100 with ResNet-50.
Table 4: ToW for different unlearning algorithms applied to forget / retain sets with varying ES, evaluated on CIFAR-10 using ResNet-18 and CIFAR-100 using ResNet-50. Across all algorithms, including the baseline without any unlearning performed (denoted as “Original"), we observe that ToW decreases from low to high ES partitions, indicating that unlearning becomes harder as the forget and retain sets become more entangled.
@@
In Section 4.2, we discussed another critical factor affecting unlearning difficulty: the memorization level of the forget examples. Specifically, when a forget set consists of examples that are more highly memorized by the model, it becomes more difficult to unlearn (for most algorithms). The detailed results that support this observation are presented in Table 7, Table 8, Table 9, and Figure 8.
Table 7 illustrates the ToW for forget sets with varying memorization levels. Additionally, Figure 8 and the tables for CIFAR-10 (Table 8) and CIFAR-100 (Table 9) provide extensive data on forget, retain, and test accuracy, as well as MIA performance. For all the experiments, the average results are reported with 95% confidence intervals over 6 runs for relabelling-based algorithms and 3 runs for others.
@@
We present the details of our RUM experiment in this section. Table 10 presents the distribution of examples in the forget set for each class in the RUM experiments for both CIFAR-10 and CIFAR-100 datasets. The size of the forget set is 3000 in all the experiments. For CIFAR-10, the table lists the number of forget set examples for each class. For CIFAR-100, the table arranges the number of forget set examples in a 20x5 format for better readability. As shown in Table 10, the forget set covers examples from all classes in both CIFAR-10 and CIFAR-100 experiments.
@@
Tables 11 provide detailed results when applying RUM to different unlearning algorithms for CIFAR-10 using ResNet-18 and CIFAR-100 using ResNet-50. We selected the top-performing algorithms from previous experiments (i.e., Fine-tune, L1-sparse, NegGrad+, SalUn) and applied the refinement strategy to each algorithm in three ways: i) “vanilla": unlearning $\mathcal{S}$ in one go, ii)“shuffle": sequentially applying the algorithm on 3 equal-sized random subsets, serving as a control experiment, iii)“RUMF": sequentially applying the algorithm on 3 subsets of $\mathcal{S}$ in low $\rightarrow$ med $\rightarrow$ high memorization order. Each experiment was conducted 3 times, with average values and 95% confidence intervals reported. Our results indicate that RUMF improves the performance of each unlearning algorithm, and the full RUM approach, which uses the best algorithm for each subset, achieves the overall best results.
@@
Furthermore, to gain a deeper understanding of the dynamics involved in applying RUM, we plotted the sequence dynamics for SalUn RUMF on CIFAR-10 (Figure 4) and NegGrad+ RUMF on CIFAR-100 (Figure 9). These plots show the accuracy of the entire forget set, the retain set, the test set, and subsets of the forget set after each step. They demonstrate that while both orderings (low $\rightarrow$ med $\rightarrow$ high and high $\rightarrow$ med $\rightarrow$ low memorization) yield similar ToW according to Table 11, their sequence dynamics during the unlearning phase are different. This phenomenon is discussed in Section 6 in the main paper, specifically in the Analysis of Sequence Dynamics paragraph.

### v2

@@ A.8 Detailed results
The first factor affecting unlearning difficulty, as discussed in Section 4.1, is the degree of entanglement between the forget and retain sets in the embedding space. Specifically, the more entangled these sets are, the harder it becomes to unlearn. We present detailed results on how this factor affects unlearning difficulty in Table 6(d), 7(c), 8(c), 9(c), 10(c) and Figure 11. Table 6(d) displays the ToW results for different ES partitions across various settings, including CIFAR-10, CIFAR-100 and Tiny-ImageNet datasets, as well as ResNet and VGG architectures. Additionally, comprehensive data—covering forget, retain, and test accuracy along with MIA performance—are provided in Figure 11 and in Table 7(c) through Table 10(c) for these datasets and model architectures. All the results are averaged over 3 runs for each algorithm (6 runs for relabelling-based algorithms due to their higher variance), along with 95% confidence intervals.
Table 6: ToW for different unlearning algorithms applied to forget / retain sets with varying ES, evaluated on various datasets and model architectures. Across all algorithms, including the baseline without any unlearning performed (denoted as “Original”), we observe that ToW decreases from low to high ES partitions, indicating that unlearning becomes harder as the forget and retain sets become more entangled.
@@
(a) CIFAR-10 with ResNet-18
@@
(b) CIFAR-100 with ResNet-50
Low ES Medium ES High ES
Retrain 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000
Original 0.920 $\pm$ 0.019 0.626 $\pm$ 0.046 0.377 $\pm$ 0.015
Fine-tune 0.919 $\pm$ 0.015 0.770 $\pm$ 0.039 0.765 $\pm$ 0.022
L1-sparse 0.920 $\pm$ 0.018 0.884 $\pm$ 0.014 0.866 $\pm$ 0.004
NegGrad 0.895 $\pm$ 0.032 0.616 $\pm$ 0.070 0.476 $\pm$ 0.027
NegGrad+ 0.921 $\pm$ 0.021 0.914 $\pm$ 0.041 0.854 $\pm$ 0.021
Influence Unlearning 0.892 $\pm$ 0.029 0.595 $\pm$ 0.019 0.442 $\pm$ 0.026
SalUn 0.920 $\pm$ 0.018 0.613 $\pm$ 0.044 0.579 $\pm$ 0.021
(c) Tiny-ImageNet with ResNet-18
Low ES Medium ES High ES
Retrain 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000
Original 0.933 $\pm$ 0.002 0.568 $\pm$ 0.017 0.331 $\pm$ 0.030
Fine-tune 0.936 $\pm$ 0.003 0.749 $\pm$ 0.055 0.677 $\pm$ 0.087
L1-sparse 0.936 $\pm$ 0.006 0.763 $\pm$ 0.043 0.737 $\pm$ 0.026
NegGrad 0.919 $\pm$ 0.019 0.569 $\pm$ 0.037 0.379 $\pm$ 0.016
NegGrad+ 0.934 $\pm$ 0.015 0.806 $\pm$ 0.027 0.577 $\pm$ 0.181
Influence Unlearning 0.933 $\pm$ 0.005 0.547 $\pm$ 0.041 0.423 $\pm$ 0.046
SalUn 0.934 $\pm$ 0.001 0.554 $\pm$ 0.010 0.411 $\pm$ 0.118
(d) Tiny-ImageNet with VGG-16
@@
Table 9: Accuracy and MIA performance of different unlearning algorithms on forget / retain sets with varying ES for Tiny-ImageNet using ResNet-18.
Forget Acc Retain Acc Test Acc MIA
Retrain 92.175 $\pm$ 1.804 99.979 $\pm$ 0.001 64.593 $\pm$ 1.146 0.164 $\pm$ 0.016
Original 100.000 $\pm$ 0.000 99.979 $\pm$ 0.000 64.773 $\pm$ 1.027 0.001 $\pm$ 0.001
Fine-tune 100.000 $\pm$ 0.000 99.979 $\pm$ 0.000 64.253 $\pm$ 0.553 0.007 $\pm$ 0.003
L1-sparse 100.000 $\pm$ 0.000 99.979 $\pm$ 0.000 64.393 $\pm$ 0.596 0.001 $\pm$ 0.002
NegGrad 99.600 $\pm$ 0.166 98.858 $\pm$ 0.319 62.386 $\pm$ 1.957 0.011 $\pm$ 0.004
NegGrad+ 100.000 $\pm$ 0.000 99.979 $\pm$ 0.000 64.506 $\pm$ 1.275 $0$$\pm$ 0.000
Influence unlearning 99.789 $\pm$ 0.048 99.637 $\pm$ 0.492 61.452 $\pm$ 2.375 0.009 $\pm$ 0.006
SalUn 100.000 $\pm$ 0.000 99.979 $\pm$ 0.000 64.513 $\pm$ 1.446 $0$$\pm$ 0.000
(a) Low ES
Forget Acc Retain Acc Test Acc MIA
Retrain 62.800 $\pm$ 4.303 99.980 $\pm$ 0.003 64.480 $\pm$ 0.698 0.670 $\pm$ 0.031
Original 99.978 $\pm$ 0.048 99.980 $\pm$ 0.001 64.773 $\pm$ 1.027 0.010 $\pm$ 0.009
Fine-tune 85.256 $\pm$ 2.155 99.980 $\pm$ 0.001 63.893 $\pm$ 1.599 0.656 $\pm$ 0.010
L1-sparse 64.989 $\pm$ 7.874 97.985 $\pm$ 1.636 56.718 $\pm$ 2.988 0.502 $\pm$ 0.015
NegGrad 89.789 $\pm$ 9.586 90.936 $\pm$ 10.713 57.258 $\pm$ 7.332 0.165 $\pm$ 0.074
NegGrad+ 68.878 $\pm$ 2.883 99.977 $\pm$ 0.007 61.792 $\pm$ 0.516 0.495 $\pm$ 0.049
Influence unlearning 82.744 $\pm$ 2.567 86.616 $\pm$ 3.098 50.297 $\pm$ 3.316 0.145 $\pm$ 0.102
SalUn 87.344 $\pm$ 3.792 86.758 $\pm$ 4.405 58.045 $\pm$ 0.977 0.689 $\pm$ 0.166
(b) Medium ES
Forget Acc Retain Acc Test Acc MIA
Retrain 37.822 $\pm$ 1.112 99.986 $\pm$ 0.006 64.313 $\pm$ 0.553 0.891 $\pm$ 0.006
Original 99.889 $\pm$ 0.172 99.983 $\pm$ 0.005 64.773 $\pm$ 1.027 0.061 $\pm$ 0.041
Fine-tune 60.889 $\pm$ 1.574 99.983 $\pm$ 0.003 63.799 $\pm$ 1.499 0.934 $\pm$ 0.008
L1-sparse 45.811 $\pm$ 0.912 99.260 $\pm$ 0.538 59.132 $\pm$ 0.849 0.813 $\pm$ 0.076
NegGrad 67.989 $\pm$ 8.858 80.444 $\pm$ 8.129 49.230 $\pm$ 4.824 0.457 $\pm$ 0.101
NegGrad+ 49.933 $\pm$ 2.191 99.982 $\pm$ 0.006 61.432 $\pm$ 0.928 0.843 $\pm$ 0.025
Influence unlearning 78.678 $\pm$ 13.145 86.881 $\pm$ 11.277 50.703 $\pm$ 5.029 0.640 $\pm$ 0.091
SalUn 62.000 $\pm$ 6.416 82.350 $\pm$ 2.534 57.091 $\pm$ 2.965 0.926 $\pm$ 0.056
(c) High ES
Table 10: Accuracy and MIA performance of different unlearning algorithms on forget / retain sets with varying ES for Tiny-ImageNet using VGG-16.
Forget Acc Retain Acc Test Acc MIA
Retrain 93.942 $\pm$ 0.752 99.979 $\pm$ 0.001 60.279 $\pm$ 1.269 0.119 $\pm$ 0.017
Original 100.000 $\pm$ 0.000 99.979 $\pm$ 0.001 60.479 $\pm$ 0.938 0.001 $\pm$ 0.003
Fine-tune 100.000 $\pm$ 0.000 99.979 $\pm$ 0.000 60.332 $\pm$ 0.398 0.001 $\pm$ 0.001
L1-sparse 100.000 $\pm$ 0.000 99.979 $\pm$ 0.000 60.225 $\pm$ 0.331 0.001 $\pm$ 0.000
NegGrad 99.978 $\pm$ 0.048 99.653 $\pm$ 0.216 58.445 $\pm$ 0.451 0.003 $\pm$ 0.003
NegGrad+ 100.000 $\pm$ 0.000 99.979 $\pm$ 0.001 59.899 $\pm$ 0.748 0.001 $\pm$ 0.002
Influence unlearning 99.989 $\pm$ 0.048 99.979 $\pm$ 0.001 60.412 $\pm$ 1.172 0.001 $\pm$ 0.001
SalUn 100.000 $\pm$ 0.000 99.979 $\pm$ 0.000 60.372 $\pm$ 1.093 0.001 $\pm$ 0.000
(a) Low ES
Forget Acc Retain Acc Test Acc MIA
Retrain 56.978 $\pm$ 2.110 99.980 $\pm$ 0.003 60.139 $\pm$ 0.445 0.662 $\pm$ 0.043
Original 99.989 $\pm$ 0.048 99.979 $\pm$ 0.000 60.479 $\pm$ 0.938 0.008 $\pm$ 0.001
Fine-tune 57.133 $\pm$ 5.043 87.325 $\pm$ 4.327 47.283 $\pm$ 1.898 0.464 $\pm$ 0.033
L1-sparse 61.633 $\pm$ 3.340 89.064 $\pm$ 0.687 49.983 $\pm$ 0.875 0.436 $\pm$ 0.049
NegGrad 88.200 $\pm$ 20.298 90.202 $\pm$ 20.598 52.604 $\pm$ 12.745 0.167 $\pm$ 0.189
NegGrad+ 69.944 $\pm$ 4.170 99.367 $\pm$ 0.642 53.284 $\pm$ 1.557 0.413 $\pm$ 0.066
Influence unlearning 81.922 $\pm$ 25.731 84.474 $\pm$ 24.099 47.469 $\pm$ 11.033 0.178 $\pm$ 0.081
SalUn 78.122 $\pm$ 6.099 78.968 $\pm$ 4.343 49.116 $\pm$ 0.508 0.251 $\pm$ 0.089
(b) Medium ES
Forget Acc Retain Acc Test Acc MIA
Retrain 33.289 $\pm$ 2.942 99.982 $\pm$ 0.004 59.859 $\pm$ 0.959 0.880 $\pm$ 0.039
Original 99.933 $\pm$ 0.083 99.981 $\pm$ 0.001 60.479 $\pm$ 0.938 0.034 $\pm$ 0.010
Fine-tune 30.089 $\pm$ 5.122 81.706 $\pm$ 5.794 45.336 $\pm$ 2.339 0.709 $\pm$ 0.055
L1-sparse 40.300 $\pm$ 3.354 88.805 $\pm$ 2.483 49.130 $\pm$ 0.358 0.688 $\pm$ 0.053
NegGrad 93.256 $\pm$ 7.453 97.402 $\pm$ 4.942 57.151 $\pm$ 2.957 0.169 $\pm$ 0.124
NegGrad+ 73.489 $\pm$ 23.035 99.802 $\pm$ 0.736 56.711 $\pm$ 3.155 0.513 $\pm$ 0.406
Influence unlearning 82.933 $\pm$ 4.683 92.497 $\pm$ 1.794 50.690 $\pm$ 0.650 0.233 $\pm$ 0.249
SalUn 84.089 $\pm$ 23.491 91.948 $\pm$ 13.228 51.810 $\pm$ 9.754 0.334 $\pm$ 0.193
(c) High ES
@@
In Section 4.2, we discussed another critical factor affecting unlearning difficulty: the memorization level of the forget examples. Specifically, when a forget set consists of examples that are more highly memorized by the model, it becomes more difficult to unlearn (for most algorithms). The detailed results that support this observation are presented in Table 11(b), Table 12(c), Table 13(c), and Figure 12.
Table 11(b) illustrates the ToW for forget sets with varying memorization levels. Additionally, Figure 12 and Table 12(c) to Table 13(c) provide extensive data on forget, retain, and test accuracy, as well as MIA performance. For all the experiments, the average results are reported with 95% confidence intervals over 6 runs for relabelling-based algorithms and 3 runs for others.
@@
Table 14: Unlearning performance evaluated by ToW-MIA, with ToW results included for comparison. Subtables (a, b): ToW-MIA for different unlearning algorithms applied to forget / retain sets with varying ES or memorization levels on CIFAR-10 using ResNet-18, with ToW results from Tables 6(d) and 11(b) for comparison. Subtable (c): RUMF and RUM results evaluated by ToW-MIA on CIFAR-10 using ResNet-18, with ToW results from Table 16(b) for comparison.
ToW-MIA ToW
Low ES Medium ES High ES Low ES Medium ES High ES
Retrain 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000
Original 0.893 $\pm$ 0.014 0.841 $\pm$ 0.026 0.430 $\pm$ 0.035 0.944 $\pm$ 0.014 0.928 $\pm$ 0.022 0.759 $\pm$ 0.048
Fine-tune 0.921 $\pm$ 0.061 0.882 $\pm$ 0.020 0.768 $\pm$ 0.055 0.952 $\pm$ 0.024 0.923 $\pm$ 0.019 0.908 $\pm$ 0.018
L1-sparse 0.895 $\pm$ 0.014 0.845 $\pm$ 0.031 0.706 $\pm$ 0.103 0.945 $\pm$ 0.013 0.926 $\pm$ 0.019 0.836 $\pm$ 0.031
NegGrad 0.883 $\pm$ 0.030 0.841 $\pm$ 0.044 0.479 $\pm$ 0.010 0.929 $\pm$ 0.030 0.887 $\pm$ 0.047 0.766 $\pm$ 0.042
NegGrad+ 0.893 $\pm$ 0.026 0.877 $\pm$ 0.057 0.550 $\pm$ 0.055 0.941 $\pm$ 0.029 0.920 $\pm$ 0.038 0.800 $\pm$ 0.029
SCRUB 0.895 $\pm$ 0.014 0.884 $\pm$ 0.045 0.778 $\pm$ 0.075 0.944 $\pm$ 0.014 0.939 $\pm$ 0.021 0.920 $\pm$ 0.033
Influence unlearning 0.879 $\pm$ 0.022 0.812 $\pm$ 0.020 0.618 $\pm$ 0.010 0.928 $\pm$ 0.023 0.890 $\pm$ 0.012 0.744 $\pm$ 0.029
SalUn 0.339 $\pm$ 0.137 0.359 $\pm$ 0.178 0.597 $\pm$ 0.126 0.783 $\pm$ 0.031 0.656 $\pm$ 0.046 0.618 $\pm$ 0.056
Random-label 0.171 $\pm$ 0.037 0.446 $\pm$ 0.219 0.557 $\pm$ 0.086 0.767 $\pm$ 0.107 0.663 $\pm$ 0.055 0.443 $\pm$ 0.087
(a) ToW-MIA vs ES
ToW-MIA ToW
Low mem Medium mem High mem Low mem Medium mem High mem
Retrain 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000 1.000 $\pm$ 0.000
Original 0.973 $\pm$ 0.037 0.340 $\pm$ 0.072 0.041 $\pm$ 0.017 0.988 $\pm$ 0.007 0.723 $\pm$ 0.053 0.231 $\pm$ 0.058
Fine-tune 0.935 $\pm$ 0.032 0.842 $\pm$ 0.037 0.738 $\pm$ 0.068 0.933 $\pm$ 0.052 0.884 $\pm$ 0.019 0.760 $\pm$ 0.065
L1-sparse 0.904 $\pm$ 0.098 0.700 $\pm$ 0.018 0.642 $\pm$ 0.025 0.914 $\pm$ 0.061 0.816 $\pm$ 0.011 0.629 $\pm$ 0.087
NegGrad 0.933 $\pm$ 0.054 0.466 $\pm$ 0.024 0.258 $\pm$ 0.105 0.938 $\pm$ 0.028 0.738 $\pm$ 0.005 0.325 $\pm$ 0.098
NegGrad+ 0.973 $\pm$ 0.028 0.523 $\pm$ 0.037 0.600 $\pm$ 0.123 0.965 $\pm$ 0.020 0.831 $\pm$ 0.032 0.661 $\pm$ 0.082
SCRUB 0.979 $\pm$ 0.011 0.913 $\pm$ 0.158 0.865 $\pm$ 0.050 0.988 $\pm$ 0.010 0.923 $\pm$ 0.033 0.780 $\pm$ 0.101
Influence unlearning 0.973 $\pm$ 0.030 0.507 $\pm$ 0.046 0.357 $\pm$ 0.061 0.986 $\pm$ 0.031 0.738 $\pm$ 0.051 0.381 $\pm$ 0.037
SalUn 0.054 $\pm$ 0.030 0.674 $\pm$ 0.097 0.963 $\pm$ 0.024 0.774 $\pm$ 0.043 0.758 $\pm$ 0.053 0.886 $\pm$ 0.082
Random-label 0.328 $\pm$ 0.394 0.722 $\pm$ 0.119 0.862 $\pm$ 0.084 0.709 $\pm$ 0.029 0.548 $\pm$ 0.093 0.877 $\pm$ 0.064
(b) ToW-MIA vs memorization
ToW-MIA ToW
Retrain 1.000 $\pm$ 0.000 1.000 $\pm$ 1.000
Fine-tune 0.820 $\pm$ 0.042 0.849 $\pm$ 0.030
Fine-tune shuffle 0.848 $\pm$ 0.125 0.712 $\pm$ 0.040
Fine-tune RUMF 0.888 $\pm$ 0.022 0.937 $\pm$ 0.052
SalUn 0.602 $\pm$ 0.019 0.731 $\pm$ 0.070
SalUn shuffle 0.677 $\pm$ 0.023 0.727 $\pm$ 0.030
SalUn RUMF 0.878 $\pm$ 0.049 0.887 $\pm$ 0.069
NegGrad+ 0.707 $\pm$ 0.033 0.802 $\pm$ 0.028
NegGrad+ shuffle 0.470 $\pm$ 0.027 0.632 $\pm$ 0.022
NegGrad+ RUMF 0.779 $\pm$ 0.025 0.879 $\pm$ 0.068
L1-sparse 0.772 $\pm$ 0.014 0.794 $\pm$ 0.035
L1-sparse shuffle 0.624 $\pm$ 0.039 0.716 $\pm$ 0.023
L1-sparse RUMF 0.855 $\pm$ 0.023 0.900 $\pm$ 0.020
[…]
(c) ToW-MIA vs RUM
@@
We present the details of our RUM experiment in this section. Table 15(b) presents the distribution of examples in the forget set for each class in the RUM experiments for both CIFAR-10 and CIFAR-100 datasets. The size of the forget set is 3000 in all the experiments. For CIFAR-10, the table lists the number of forget set examples for each class. For CIFAR-100, the table arranges the number of forget set examples in a 20x5 format for better readability. As shown in Table 15(b), the forget set covers examples from all classes in both CIFAR-10 and CIFAR-100 experiments.
@@
Tables 16(b) provide detailed results when applying RUM to different unlearning algorithms for CIFAR-10 using ResNet-18 and CIFAR-100 using ResNet-50. We selected the top-performing algorithms from previous experiments (i.e., Fine-tune, L1-sparse, NegGrad+, SalUn) and applied the refinement strategy to each algorithm in three ways: i) “vanilla": unlearning $\mathcal{S}$ in one go, ii)“shuffle": sequentially applying the algorithm on 3 equal-sized random subsets, serving as a control experiment, iii)“RUMF": sequentially applying the algorithm on 3 subsets of $\mathcal{S}$ in low $\rightarrow$ med $\rightarrow$ high memorization order. Each experiment was conducted 3 times, with average values and 95% confidence intervals reported. Our results indicate that RUMF improves the performance of each unlearning algorithm, and the full RUM approach, which uses the best algorithm for each subset, achieves the overall best results.
@@
Furthermore, to gain a deeper understanding of the dynamics involved in applying RUM, we plotted the sequence dynamics for SalUn RUMF on CIFAR-10 (Figure 5) and NegGrad+ RUMF on CIFAR-100 (Figure 13). These plots show the accuracy of the entire forget set, the retain set, the test set, and subsets of the forget set after each step. They demonstrate that while both orderings (low $\rightarrow$ med $\rightarrow$ high and high $\rightarrow$ med $\rightarrow$ low memorization) yield similar ToW according to Table 16(b), their sequence dynamics during the unlearning phase are different. This phenomenon is discussed in Section 6 in the main paper, specifically in the Analysis of Sequence Dynamics paragraph.
@@
Table 17: RUM results using C-proxy for the refinement step, evaluated by ToW on CIFAR-10 and Tiny-ImageNet datasets. Each algorithm is applied in three ways: RUMF, as well as vanilla and shuffle as comparison. Results are averaged over 3 runs with 95% confidence intervals.
Retrain 1.000$\pm$0.000 50.433$\pm$6.808 100.000$\pm$0.000 84.167$\pm$1.616 0.637$\pm$0.032 0.000
Fine-tune 0.731$\pm$0.021 77.389$\pm$3.298 97.443$\pm$2.263 81.847$\pm$1.021 0.286$\pm$0.062 0.351
Fine-tune shuffle 0.829$\pm$0.022 62.800$\pm$16.311 98.060$\pm$5.309 80.677$\pm$5.082 0.560$\pm$0.091 0.077
Fine-tune RUMF 0.927$\pm$0.047 58.400$\pm$1.723 98.920$\pm$2.484 82.007$\pm$3.186 0.482$\pm$0.004 0.155
L1-sparse 0.754$\pm$0.071 66.856$\pm$9.041 96.095$\pm$4.676 78.200$\pm$5.173 0.492$\pm$0.041 0.145
L1-sparse shuffle 0.618$\pm$0.095 83.222$\pm$1.391 95.890$\pm$4.480 79.963$\pm$5.127 0.256$\pm$0.031 0.381
L1-sparse RUMF 0.907$\pm$0.026 53.611$\pm$3.446 96.947$\pm$1.991 80.783$\pm$2.540 0.566$\pm$0.021 0.071
NegGrad+ 0.880$\pm$0.039 56.067$\pm$5.261 99.037$\pm$1.532 78.443$\pm$5.531 0.504$\pm$0.049 0.133
NegGrad+ shuffle 0.529$\pm$0.071 94.867$\pm$9.081 98.919$\pm$2.120 80.563$\pm$6.630 0.110$\pm$0.107 0.527
NegGrad+ RUMF 0.893$\pm$0.026 58.400$\pm$1.723 98.920$\pm$2.484 82.007$\pm$3.186 0.482$\pm$0.004 0.155
SalUn 0.859$\pm$0.042 58.111$\pm$4.816 98.030$\pm$1.356 79.180$\pm$4.306 0.593$\pm$0.014 0.044
SalUn shuffle 0.638$\pm$0.071 84.067$\pm$5.039 97.897$\pm$0.449 82.387$\pm$3.331 0.381$\pm$0.015 0.256
SalUn RUMF 0.737$\pm$0.040 75.578$\pm$4.606 99.996$\pm$0.009 82.597$\pm$4.474 0.949$\pm$0.049 0.312
SCRUB vanilla 0.802$\pm$0.047 71.733$\pm$2.630 99.671$\pm$0.708 82.077$\pm$4.319 0.462$\pm$0.031 0.175
SCRUB shuffle 0.732$\pm$0.059 74.678$\pm$20.089 95.617$\pm$11.622 80.983$\pm$8.000 0.325$\pm$0.163 0.312
SCRUB RUMF 0.945$\pm$0.017 50.589$\pm$0.526 99.375$\pm$0.560 82.487$\pm$3.277 0.500$\pm$0.003 0.137
(a) ToW vs RUM on CIFAR-10 with ResNet-18
Retrain 1.000$\pm$0.000 45.633$\pm$2.521 99.999$\pm$0.001 64.606$\pm$0.230 0.662$\pm$0.005 0.000
Fine-tune 0.777$\pm$0.020 52.356$\pm$3.990 91.768$\pm$2.396 55.438$\pm$1.704 0.530$\pm$0.024 0.132
Fine-tune shuffle 0.566$\pm$0.021 80.622$\pm$0.981 94.195$\pm$0.645 56.971$\pm$1.295 0.276$\pm$0.026 0.386
Fine-tune RUMF 0.803$\pm$0.023 51.578$\pm$3.275 93.174$\pm$1.606 56.185$\pm$1.393 0.554$\pm$0.032 0.108
L1-sparse 0.772$\pm$0.017 51.622$\pm$2.343 89.556$\pm$1.293 56.258$\pm$0.965 0.534$\pm$0.025 0.128
L1-sparse shuffle 0.584$\pm$0.025 82.478$\pm$0.172 98.426$\pm$0.629 58.498$\pm$1.567 0.244$\pm$0.035 0.418
L1-sparse RUMF 0.883$\pm$0.032 55.722$\pm$0.485 99.985$\pm$0.058 62.833$\pm$0.258 0.660$\pm$0.007 0.002
NegGrad+ 0.859$\pm$0.066 51.122$\pm$11.593 97.359$\pm$1.060 57.992$\pm$1.441 0.677$\pm$0.160 0.015
NegGrad+ shuffle 0.482$\pm$0.004 85.433$\pm$21.932 93.073$\pm$17.340 51.710$\pm$18.821 0.568$\pm$0.648 0.094
NegGrad+ RUMF 0.935$\pm$0.040 46.944$\pm$4.332 99.666$\pm$0.056 59.952$\pm$1.194 0.621$\pm$0.019 0.041
SalUn 0.689$\pm$0.008 67.178$\pm$2.079 89.889$\pm$0.888 62.319$\pm$0.748 0.556$\pm$0.209 0.106
SalUn shuffle 0.553$\pm$0.032 65.333$\pm$5.584 80.933$\pm$7.695 49.750$\pm$3.141 0.397$\pm$0.030 0.265
SalUn RUMF 0.736$\pm$0.027 54.867$\pm$3.589 86.485$\pm$2.309 63.151$\pm$2.853 0.560$\pm$0.024 0.102
(b) ToW vs RUM on Tiny-ImageNet with ResNet-18
Retrain 1.000$\pm$0.000 49.100$\pm$1.910 99.995$\pm$0.003 60.699$\pm$0.207 0.637$\pm$0.013 0.000
NegGrad+ 0.637$\pm$0.106 79.356$\pm$31.411 97.146$\pm$11.969 55.678$\pm$12.037 0.301$\pm$0.188 0.336
NegGrad+ shuffle 0.514$\pm$0.009 78.078$\pm$42.146 86.404$\pm$29.465 47.103$\pm$25.866 0.376$\pm$0.598 0.261
NegGrad+ RUMF 0.812$\pm$0.053 57.233$\pm$9.692 97.515$\pm$9.221 51.430$\pm$6.930 0.477$\pm$0.062 0.160
L1-sparse 0.771$\pm$0.030 53.167$\pm$2.384 89.907$\pm$1.114 50.117$\pm$1.047 0.500$\pm$0.031 0.137
L1-sparse shuffle 0.589$\pm$0.028 83.778$\pm$0.539 96.793$\pm$0.110 53.924$\pm$1.024 0.217$\pm$0.025 0.420
L1-sparse RUMF 0.885$\pm$0.024 56.678$\pm$1.315 99.932$\pm$0.080 56.471$\pm$2.238 0.585$\pm$0.024 0.052
Fine-tune 0.750$\pm$0.013 49.433$\pm$1.168 87.214$\pm$2.139 47.343$\pm$1.151 0.528$\pm$0.016 0.109
Fine-tune shuffle 0.576$\pm$0.061 65.978$\pm$11.740 81.317$\pm$8.350 46.023$\pm$4.367 0.360$\pm$0.160 0.277
Fine-tune RUMF 0.767$\pm$0.044 49.989$\pm$2.632 90.010$\pm$3.150 46.896$\pm$2.862 0.527$\pm$0.024 0.110
SalUn 0.602$\pm$0.041 85.522$\pm$2.569 98.336$\pm$0.254 56.945$\pm$1.357 0.322$\pm$0.035 0.315
SalUn shuffle 0.566$\pm$0.011 72.622$\pm$43.082 84.977$\pm$36.200 50.430$\pm$15.871 0.313$\pm$0.380 0.324
SalUn RUMF 0.679$\pm$0.025 54.622$\pm$5.034 82.377$\pm$2.398 47.990$\pm$1.540 0.474$\pm$0.136 0.163
(c) ToW vs RUM on Tiny-ImageNet with VGG-16